How many five-digit numbers may be made from the five digits 1, 2, 3, 4, 5.
If a digit may be used exactly once and neither of the first two digits can be 5?

Answer by ikleyn(52778)

.
How many five-digit numbers may be made from the five digits 1, 2, 3, 4, 5.
If a digit may be used exactly once and neither of the first two digits can be 5?
~~~~~~~~~~~~~~~~~~

In all, you have 5 digits in your possession.


As 1st digit of your number, you may have any of 4 digits {1,2,3,4}.           4 options.


As 2nd digit of your number, you may have any of 3 remaining digits.           3 options


As 3rd digit of your number, you can use any of remaining 5-2 = 3 digits.      3 options.


As 4th digit of your number, you can use any of remaining 5-3 = 2 digits.      2 options.


As 5th digit of your number, you can use any of remaining 5-4 = 1 digits.      1 options.


So, the answer to the problem's question is  4*3*3*2*1 = 72  possible 5-digit numbers under the imposed conditions.

Solved and carefully/thoroughly explained.